OpenStudy (anonymous):

which of the following is an identity? A. csc^2x+cot^2x=1 B. (cscx+cotx)^2=1 C. sin^2x-cos^2x=1 D. sin^2x sec^2x+1=tan^2x csc^2x

Miracrown (miracrown):

Which one do you think?

OpenStudy (anonymous):

wait

Miracrown (miracrown):

@magepker728 - Can you please refrain from giving users direct answers, as it hinders the learning process and is strongly prohibited according to the OS CoC. Thank you.

OpenStudy (anonymous):

@Miracrown back off

Miracrown (miracrown):

Excuse me? I'm trying to enforce the rules here and you're evading it. Wow, that'll get you to places for sure!

Miracrown (miracrown):

@chalkzone - we don't know in advance let's start with A and see if we can prove whether it is an identity or not let's simplify the left side, and see what it makes. convert it into terms of cos and sin csc^2x + cot^2x = 1 Can you do that?

OpenStudy (anonymous):

no not at all sadface

Miracrown (miracrown):

\[\frac{ 1 }{ \sin ^{2} x } \space + \frac{ \cos ^{2} x }{ \sin ^{2} x }\]

Miracrown (miracrown):

\[\frac{ 1+\cos ^{2}x }{ \sin ^{2}x }\] \[\frac{ 1+\cos ^{2} x}{ 1-\cos ^{2} x }\] is this an identity?

OpenStudy (anonymous):

yes! I just looked at my notes

Miracrown (miracrown):

Hmm, well, I don't think so

Miracrown (miracrown):

in fact, I am quite certain there are only a few x values for which this is true

OpenStudy (anonymous):

oh nevermind, the one I saw on my paper was vice versa

Miracrown (miracrown):

here is the deal: let's pretend it is an identity then:

Miracrown (miracrown):

\[1+\cos^2x \space = 1-\cos^2x\] \[2\cos^2x = 0\] \[cosx = 0\]

Miracrown (miracrown):

is cos x = 0 for all x?

OpenStudy (anonymous):

don't know

Miracrown (miracrown):

It should be no. Only specific x values make this true and identity is an equation that is true for ALL x... or, at least, true for all but maybe a few, like no dividing by zero, or something like that

Miracrown (miracrown):

this is not an identity. let's look at b, and apply the same strategy

Miracrown (miracrown):

\[(cscx+cotx)^2 = 1\]

Miracrown (miracrown):

again, let's convert to sin and cos, like before ...

Miracrown (miracrown):

\[(\frac{ 1 }{ sinx } \space +\frac{ cosx }{ sinx } )^2 \space = 1\] \[(\frac{ 1+cosx }{ sinx } )^2 \space = 1\] \[\frac{ 1+2cosx+\cos^2x }{ \sin^2x } \space = 1\]

Miracrown (miracrown):

\[\frac{ 1+2cosx+\cos^2x }{ 1-\cos^2x } = 1\]

Miracrown (miracrown):

Predictions?

OpenStudy (anonymous):

I am so clueless right now

Miracrown (miracrown):

\[1+2cosx+\cos^2x \space =1-\cos^2x\] \[2cosx+2\cos^2x \space = 0\] \[cosx(1+cosx) = 0\] \[cosx= 0 , -1\] Is this an identity?

Miracrown (miracrown):

Let me give a hand: Its not an identity This may seem strange, but when we find the one that is an identity, you will see the difference

Miracrown (miracrown):

\[\sin^2x-\cos^2x=1\] \[-\cos2x=1\] \[\cos2x=-1\] True for all x?

OpenStudy (anonymous):

no

Miracrown (miracrown):

Right! :) So the answer must be D, but let's see why:

Miracrown (miracrown):

\[\sin^2xsec^2x+1 = \tan^2 x \csc^2x\] \[\frac{ \sin^2x }{ \cos^2x } + 1 \space = \frac{ \sin^2x }{ \cos^2x } \space \frac{ 1 }{ \sin^2x }\] \[\tan^2x+1 = \sec^2x\] \[\sin^2x+\cos^2x=1\]

Miracrown (miracrown):

How about this?

OpenStudy (anonymous):

oooooooo damn so that's what you need to do to get the equation

Miracrown (miracrown):

Yes! So now we see that this is an identity, because, using known identities, we reduce this to an expression where the left side is IDENTICAL to the right side. so this is true, for ALL x hence why we call it an IDENTITY

OpenStudy (anonymous):

I was so confused

OpenStudy (anonymous):

thanks

Miracrown (miracrown):

yw :)